Circular No. 7715

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1650–500 M. Revnivtsev and R. Sunyaev, Max-Planck-Institut für Astrophysik

and Russian Space Research Institute, write: “Analysis of public-domain data of the target-of-opportunity RXTE observations of XTE J1650–500 (cf. IAUC 7707, 7708, 7710) from Sept. 9 and 10 reveals the presence of two quasiperiodic-oscillation (QPO) peaks (likely harmonics). The centroidpeak frequencies changed from 1.13 ± 0.02 and 2.28 ± 0.02 Hz on Sept. 9.1 UT to 1.33 ± 0.02 and 2.73 ± 0.03 Hz on Sept. 10.5. The amplitudes of the QPOs are ∼ 5–7 percent. The preliminary spectral analysis of PCA (3–20 keV) and HEXTE (20–200 keV) data shows that the source has a spectrum similar to that of Cyg X-1 and other black-hole binaries in the low/hard spectral state. The spectrum has a power-law-like shape with photon index around 1.5–1.8 and exponential cutoff with e-folding value ∼ 90 keV. The investigation of reported QPOs in the optical band could provide additional information about properties of the accretion disk and the binary.”

R. Wijnands, J. M. Miller, and W. H. G. Lewin, Massachusetts Institute of Technology, report on behalf of a larger collaboration: “A timing analysis of the RXTE PCA data obtained for the new x-ray transient XTE J1650–500 on four occasions on Sept. 6.98 (0.37 Crab), 7.13 (0.39 Crab), 8.38 (0.44 Crab), and 9.11 UT (0.45 Crab) demonstrates that the rapid x-ray variability of this source can be characterized by a strong (∼ 22 percent rms amplitude in the energy band 2–60 keV), band-limited noise component, which follows a power law with index of 1–1.5 at high frequencies with a break at 0.2 Hz, below which the spectrum is relatively flat. Superimposed on this noise component, broad (width of 0.5–1.5 Hz) features or QPOs are present with a frequency of 1.6–2.3 Hz. Also a broad (width of 0.4 Hz) feature at half the QPO frequency is detected. An example of the power-density spectrum of the source can be found

at http://space.mit.edu/ rudy/XTEJ1650-500PDS.ps (obtained for the

observation on Sept. 9.11). The spectral shape (IAUC 7707) and the timing behavior indicate that this is a black-hole binary in the low (spectrally hard) state.”

M. M. Buxton, Research School of Astronomy and Astrophysics, Australian National University; and D. Londish, University of Sydney, report that the optical counterpart of XTE J1650–500 was visible at B = 18.21 ± 0.2 on Sept. 10.42 UT on two 300-s CCD exposures taken with the 1.88-m telescope (+ Monash Cassegrain Imager) at Mount Stromlo Observatory.

2001 September 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green